Number One:
5x + 4y - z = 1
2x - 2y + z = 1
-x - y + z = 2

Number Two:
x + y + z = 0
2x + 3y +2z = -1
x - y + z = 2

Number Three:
x + 2y = 0
4x - z = 4
5y + z = -1

Answer by Edwin McCravy(20086)

You didn't specify which method to use, and there are many
different methods for solving systems of equations.  Therefore
I'll solve them any way I choose and just give the answers.
If you want them solved step by step, then please repost telling
us which of the many ways you are studying for solving these
systems. 

Number One:
5x + 4y - z = 1
2x - 2y + z = 1
-x - y + z = 2

Solution is (x,y,z,) = (0,1,3)

Number Two:
x + y + z = 0
2x + 3y +2z = -1
x - y + z = 2

Solution is (x,y,z,) = (1-a,-1,a) where a 
is an arbitrary number.


Number Three:
x + 2y     = 0
4x     - z = 4
    5y + z = -1

Solution is (x,y,z,) = (2,-1,4)

Edwin